# BRIDGE: Bootstrapping Text to Control Time-Series Generation via Multi-Agent Iterative Optimization and Diffusion Modeling

### v5

@@ F.1 Details of Datasets
Electricity: This dataset captures hourly electricity consumption for 321 clients between 2012 and 2014, measured in kilowatts (kW). It was originally sourced from the UCI repository.
Solar: Comprising 137 time series, this dataset records hourly solar power production in the state of Alabama throughout 2006.
Wind: Wind: This dataset includes a single, extensive daily time series that tracks wind power production (in megawatts) at 4-second intervals, starting from August 1, 2019. It was obtained from the Australian Energy Market Operator (AEMO) platform.
Traffic: Covering 15 months of daily data (440 records), this dataset represents the occupancy rate (ranging from 0 to 1) of various car lanes on the San Francisco Bay Area freeways over time.
Taxi This dataset contains spatio-temporal traffic time series of New York City taxi rides, recorded every 30 minutes at 1,214 locations during January 2015 and January 2016.
Pedestrian: Featuring hourly pedestrian counts from 66 sensors in Melbourne, this dataset spans from May 2009 to April 30, 2020, and is regularly updated as new data becomes available.
Air Quality: Used in the KDD Cup 2018 forecasting competition, this dataset includes hourly air quality measurements from 59 stations in Beijing (35 stations) and London (24 stations) between January 1, 2017, and March 31, 2018. The data includes various air quality metrics such as PM2.5, PM10, NO2, CO, O3, and SO2. Missing values were imputed using leading zeros or the Last Observation Carried Forward (LOCF) method.
Temperature: This dataset consists of 32,072 daily time series with temperature observations and rain forecasts from 422 weather stations across Australia, collected between May 2, 2015, and April 26, 2017. Missing values were replaced with zeros, and the mean temperature column was extracted for use.
Rain: Dataset focuses on rain data extracted from the same source as temperature.
NN5: This dataset contains 111 time series from the banking sector, with the goal of predicting daily cash withdrawals from ATMs in the UK.
Fred-MD: This dataset contains 107 monthly time series reflecting various macroeconomic indicators, sourced from the factuality Reserve Bank’s FRED-MD database. The series have been differenced and log-transformed following established practices in the literature.
@@ F.2 Dataset Statistics
Short-term Forecasting M4-Yearly 1 6 (23000, 0, 23000) Yearly Demographic
@@ H.1 Evaluation Process:
Preliminary Setup Annotators were provided with one generated text description with different number of time series. Each description was assessed independently, with annotators reviewing how well the time series represented the trends, patterns, and anomalies observed in the corresponding text descriptions, providing a ranking of time series. We provide two different settings: HE represent only ranking the 4 different settings output with ground truth time series and HE@3 represent ranking 4 different settings’ time series, ground truth time series and 3 random select time series from test set.
@@ Appendix I Performance of downstream tasks
Table 8: Comparison of MSE and MAE for TS forecasting across methods, with results for four forecasting horizons: $H$in $\{24,36,48,60\}$ for ILI and $H$in $\{6,48\}$ for M4. Average results are reported.
@@
Marginal Distribution Distance Electricity $0.110$ $0.139$ $0.135$ $0.256$
@@
Marginal Distribution Distance Electricity $0.615$ $0.368$ $0.135$ $0.117$ $0.236$ K-L Divergence $0.006$ $0.027$ $0.001$ $0.001$ $0.020$
@@
Table 12: Comparison of MSE and MAE across various methods on Long-term forecasting. The results are for four different forecasting horizons: $H$in $\{24,36,48,60\}$. Red values indicate the best score, and blue values represent the second best.

### v6

@@ F.1 Details of Datasets
Electricity: This dataset represents the hourly electricity consumption of 321 clients from 2012 to 2014, measured in kilowatts (kW). It was originally extracted from the UCI repository.
Solar: This dataset contains 137 time series representing hourly solar power production in the state of Alabama throughout 2006.
Wind: This dataset contains a single extensive daily time series representing wind power production (in megawatts) recorded at 4-second intervals starting from August 1, 2019. It was downloaded from the Australian Energy Market Operator (AEMO) online platform.
Traffic: This dataset contains 15 months of daily data (440 daily records) describing the occupancy rate (between 0 and 1) of different car lanes on San Francisco Bay Area freeways over time.
Taxi: This dataset contains spatio-temporal traffic time series of New York City taxi rides recorded at 1,214 locations every 30 minutes during January 2015 and January 2016.
Pedestrian: This dataset contains hourly pedestrian counts captured from 66 sensors in Melbourne starting from May 2009. The original dataset is regularly updated when new observations become available. The dataset used here contains pedestrian counts up to April 30, 2020.
Air Quality: This dataset was used in the KDD Cup 2018 forecasting competition. It contains hourly air quality measurements from 59 stations in two cities: Beijing (35 stations) and London (24 stations) from January 1, 2017, to March 31, 2018. The air quality measurements include various metrics such as PM2.5, PM10, NO2, CO, O3, and SO2. Missing values were imputed using leading zeros or the Last Observation Carried Forward (LOCF) method.
Temperature: This dataset contains 32,072 daily time series with temperature observations and rain forecasts gathered by the Australian Bureau of Meteorology from 422 weather stations across Australia between May 2, 2015, and April 26, 2017. Missing values were replaced with zeros, and the mean temperature column was extracted for use.
Rain: This dataset focuses on rain data extracted from the same source as the Temperature dataset.
NN5: This dataset was used in the NN5 forecasting competition. It contains 111 time series from the banking domain with the goal of predicting daily cash withdrawals from ATMs in the UK. Missing values were replaced by the median across all corresponding days of the week throughout the entire series.
Fred-MD: This dataset contains 107 monthly time series reflecting various macroeconomic indicators sourced from the Federal Reserve Bank’s FRED-MD database. The series have been differenced and log-transformed following established practices in the literature.
@@
Web: This dataset was used in the Kaggle Wikipedia Web Traffic forecasting competition. It contains 145,063 daily time series representing the number of hits or web traffic for Wikipedia pages from July 1, 2015, to September 10, 2017. Missing values were replaced with zeros.
@@ F.2 Dataset Statistics
Short-term
Forecasting
M4-Yearly 1 6 (23000, 0, 23000) Yearly Demographic
@@ H.1 Evaluation Process:
Preliminary Setup Annotators were presented with a single text description accompanied by a set of time series. Each set was evaluated independently, with annotators assessing how well each time series captured the trends, patterns, and anomalies described in the corresponding text. We define two evaluation settings: (i) HE-Rank, annotators ranked the time series generated by different model settings together with the ground truth, focusing purely on relevance. (ii) HE-Mixed, the same sequences were evaluated alongside three distractor time series randomly sampled from the test set, forming a mixed pool of candidates for relative ranking.
@@ Appendix I Performance of downstream tasks
Table 8: Comparison of MSE and MAE for TS forecasting across methods, with results for four forecasting horizons: $H$ $\in$ $\{24,36,48,60\}$ for ILI and $H$ $\in$ $\{6,48\}$ for M4. Average results are reported.
@@
Marginal Distribution Distance
Electricity $0.110$ $0.139$ $0.135$ $0.256$
@@
Marginal Distribution Distance
Electricity $0.615$ $0.368$ $0.135$ $0.117$ $0.236$
K-L Divergence
$0.006$ $0.027$ $0.001$ $0.001$ $0.020$
@@
Table 12: Comparison of MSE and MAE across various methods on Long-term forecasting. The results are for four different forecasting horizons: $H$ $\in$ $\{24,36,48,60\}$. Red values indicate the best score, and blue values represent the second best.